OpenStudy (anonymous):

What is the area of the parallelogram below? 18 21 36 42

OpenStudy (anonymous):



Directrix (directrix):

The area of a parallelogram is base times corresponding altitude (drawn to that base). A = b*h A = 12*3 = 36

OpenStudy (anonymous):

thanks!! i think

Directrix (directrix):

I assumed that the sides of the parallelogram are 3 and 14. What's unusual here is that we usually take the horizontal view on base and vertical view on height. Set the parallelogram on its short side, and you have an equally good base and height drawn to that base.